Let H(t)= 4 + 4sin(pit/2). Approximate to the nearest hundreth, all values t in the interval [0,4pi] for which H(t)=6

Answer by stanbon(75887)

Let H(t)= 4 + 4sin(pit/2). Approximate to the nearest hundreth, all values t in the interval [0,4pi] for which H(t)=6
---------------
4 + 4 sin[(pi/2)t] = 6
sin[(pi/2)t] = 1/2
(pi/2)t = (1/6)pi+2npi or (5/6)pi+2npi, n=0,1,2
t=(1/3)+4n or t=(5/3)+4n, n=0,1,2,
t= 1/3, 5/3, 13/3, 17/3, 25/3, 29/3
===========
Cheers,
Stan H.